OpenStudy (anonymous):

Five people can mow a lawn in 12 hours. How many more people are needed to mow the lawn in just 3 hours, assuming each person mows at the same rate?

OpenStudy (anonymous):

thhe answer is 15

OpenStudy (mathmate):

Five can mow in 12 hours, so one can mow in 5*12=60 hours.To mow in 3 hours, you'll need 60/3=20 people, = 15 MORE people.

OpenStudy (anonymous):

miss pac girl

OpenStudy (anonymous):

ya, I was thinking 20 people but accidently forgot to subtract the number of ppl. 20-5=15